OpenStudy (anonymous):

y=-(x+2)(x-4) find the domian,range,axis of symmetry

OpenStudy (anonymous):

@welshfella

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first step would be to multiply this out

OpenStudy (anonymous):

Okay

OpenStudy (misty1212):

let mt know what you get

OpenStudy (misty1212):

then we can finish quickly

OpenStudy (anonymous):

Alright, hold on:)

OpenStudy (anonymous):

-x^2+2x+8

OpenStudy (misty1212):

ok good

OpenStudy (anonymous):

:)

OpenStudy (misty1212):

now the fact that the leading coefficient it negative ( it is \(-1\)) means this parabola opens down all that is left is to find the vertex

OpenStudy (misty1212):

the first coordinate of the vertex of \(y=ax^2+bx+c\) is \[-\frac{b}{2a}\] which in your case is \[-\frac{2}{2\times (-1)}\]

OpenStudy (anonymous):

1

OpenStudy (misty1212):

right and the second coordinate is what you get for \(y\) when you replace \(x\) by \(1\)

OpenStudy (anonymous):

9

OpenStudy (misty1212):

right again

OpenStudy (misty1212):

so lets answer each question in order

OpenStudy (anonymous):

d=all real #'s right?

OpenStudy (misty1212):

Find the domian the domain of an polynomial is \(\mathbb{R}\) what you said

OpenStudy (misty1212):

the range since it is a parabola that opens down, and has a maximum at \(y=9\) the range is?

OpenStudy (anonymous):

y>=-9

OpenStudy (misty1212):

oh no

OpenStudy (anonymous):

Oh:( Ok

OpenStudy (misty1212):

not sure where you got the \(-9\) from it is \(9\) right?

OpenStudy (anonymous):

yes, whoops!

OpenStudy (misty1212):

and also the parabola opens down that means the MAXIMUM value is 9, not the minimum value

OpenStudy (misty1212):



OpenStudy (anonymous):

is maximum <=,>= or <,>

OpenStudy (misty1212):

lets forget about rules and think the possible values for \(y\) are numbers less than or equal to 9

OpenStudy (anonymous):

Alright

OpenStudy (misty1212):

if you want to write it as an inequality you can write \[y\leq 9\] or as an interval \[(-\infty, 9]\]

OpenStudy (anonymous):

okay, I see :)

OpenStudy (misty1212):

and finally, the axis of symmetry which we found ages ago is \(x=1\)

OpenStudy (misty1212):

done right?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

Thank you:)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]